OpenStudy (angela210793):

In a group of pupils 35% study math,75% chemistry and 4 pupils study both math and chemistry .Each of the pupils studies at least one of these 2 subjects....How many pupils study only math?

OpenStudy (angela210793):

r there only 10 ppl tht study math?

OpenStudy (angela210793):

hmmm i did it like this...let me write it...

OpenStudy (angela210793):

hmm 35+75=110%..so 10%of X=4--->x=40

OpenStudy (angela210793):

\[40=n(M \cup K)\]

OpenStudy (anonymous):

10 % study both... => 25 % study math only

OpenStudy (angela210793):

hmmm....no....i can't find the symbol here....let me use word...and i'll attach it :)

OpenStudy (anonymous):

n(M)=0.35s n(C)=0.75s n(MUC)=s n(M and C)=4 using the addition law we get s= 40 hence n(M) = 0.35*40 = 14

OpenStudy (angela210793):

yea tht's right but u have to find the no. of ppl who study ONLY math..so n(M/C)=14-4=10

OpenStudy (angela210793):

So i am right :) thnx :):)

OpenStudy (anonymous):

ohh. yeah . thts right. srry i didn't read the question correctly.

OpenStudy (angela210793):

^_^ Thnx :)